# Open-Sora 2.0: Training a Commercial-Level Video Generation Model in \$200k

### v1

@@ 6 System Optimization
We train our models using ColossalAI [undeft], an efficient parallel training system. Our hardware setup includes H200 GPUs, whose 141GB memory enables more effective data parallelism (DP) and allows for more aggressive selective activation checkpointing, significantly optimizing resource utilization. Additionally, we leverage PyTorch compile [undef] and Triton kernels [undefab] to accelerate training efficiency.
@@ 6.1 Parallelism Strategy
To efficiently handle high-resolution video training, we employ multiple parallelization techniques. For video autoencoders, we adapt tensor parallelism (TP) [undefai] to convolution layers by partitioning weights either the input or output channel dimensions, reducing memory consumption while preventing out-of-bound indexing for high-resolution training.
For MMDiT training, we combine Zero Redundancy Optimizer (ZeroDP) [undefaf] with Context Parallelism (CP) [undefw]. Here, video and text sequences are partitioned across GPUs along the sequence dimension, enabling each GPU to compute attention independently. This approach effectively mitigates the memory bottleneck and enhances the efficiency of attention computation, which is particularly beneficial for high-resolution videos, where attention complexity grows quadratically.
@@ 7 Performance
We further evaluate our model’s performance using VBench [undefp] in Table 13, demonstrating significant improvements from Open-Sora 1.2 to 2.0. The performance gap between Open-Sora and OpenAI’s Sora has been reduced from 4.52% to 0.69%, highlighting substantial advancements in video generation quality. Additionally, our model achieves a higher VBench score compared to CogVideoX1.5-5B and HunyuanVideo, further establishing its superiority among current open-source text-to-video models.

### v2

@@ 5.3 Motion Score
5.4 Inference-Time Scaling
Figure 13: Diffusion Inference Scaling Framework. At selected denoising steps, controlled noise injections produce multiple candidate outputs. Each candidate is evaluated using VBench metrics, and the highest-scoring variant is chosen to continue the generation process. This enables quality-aware inference without modifying the model itself.
[…]
These candidates are then evaluated using six VBench metrics [huang2024vbench]: subject_consistency, background_consistency, motion_smoothness, dynamic_degree, aesthetic_quality, imaging_quality. Depending on the application, users can assign higher priority to specific metrics—for example, favoring motion_smoothness for more natural movements or aesthetic_quality for improved visual appeal. The highest-scoring candidate is selected to continue in the denoising process, allowing the model to explore a richer generation space with stronger alignment to target attributes.
The computational overhead of this strategy depends on four primary factors: the frequency of noise injection, the number of initial noise seeds, the number of variations per step, and the number of active metric verifiers. In our experiments, we fix the verifier set to isolate the effects of the other three variables.
As shown in Figure 14, this method significantly improves generation quality under challenging prompts. The baseline (top row) exhibits static or unnatural motion, while moderate scaling (middle row) shows limited gains. In contrast, full scaling (bottom row) produces stable and natural motion sequences with improved visual coherence and dynamics.
While inference-time scaling improves overall video quality, we do not include it in our evaluation comparisons with other models to ensure a fair and consistent benchmarking setting.
Figure 14: Inference Scaling Effects. Comparison of generated bird motion across three settings. Top: Baseline shows static or unnatural motion. Middle: Partial scaling provides limited enhancement. Bottom: Full scaling produces smooth and natural motion through quality-guided inference.
@@ 6 System Optimization
We train our models using ColossalAI [10.1145/3605573.3605613], an efficient parallel training system. Our hardware setup includes H200 GPUs, whose 141GB memory enables more effective data parallelism (DP) and allows for more aggressive selective activation checkpointing, significantly optimizing resource utilization. Additionally, we leverage PyTorch compile [Ansel_PyTorch_2_Faster_2024] and Triton kernels [triton_language] to accelerate training efficiency.
@@ 6.1 Parallelism Strategy
To efficiently handle high-resolution video training, we employ multiple parallelization techniques. For video autoencoders, we adapt tensor parallelism (TP) [shoeybi2020megatronlmtrainingmultibillionparameter] to convolution layers by partitioning weights either the input or output channel dimensions, reducing memory consumption while preventing out-of-bound indexing for high-resolution training.
For MMDiT training, we combine Zero Redundancy Optimizer (ZeroDP) [rajbhandari2020zeromemoryoptimizationstraining] with Context Parallelism (CP) [liu2023ringattentionblockwisetransformers]. Here, video and text sequences are partitioned across GPUs along the sequence dimension, enabling each GPU to compute attention independently. This approach effectively mitigates the memory bottleneck and enhances the efficiency of attention computation, which is particularly beneficial for high-resolution videos, where attention complexity grows quadratically.
@@ 7 Performance
We further evaluate our model’s performance using VBench [huang2024vbench] in Figure 16, demonstrating significant improvements from Open-Sora 1.2 to 2.0. The performance gap between Open-Sora and OpenAI’s Sora has been reduced from 4.52% to 0.69%, highlighting substantial advancements in video generation quality. Additionally, our model achieves a higher VBench score compared to CogVideoX1.5-5B and HunyuanVideo, further establishing its superiority among current open-source text-to-video models.